Mr Hendricks invested R6000 at 7% p.a for 3 years in a fix deposit account. What is the interest earned over 3 years and what is the amount paid to Mr Hendricks on Maturity?


1
Expert's answer
2021-02-02T04:58:20-0500

Maturity value:

"A=P(1+i)^n"

"P" is principal amount, "i" is interest rate, "n" is the number of compounding periods

"A=6000\cdot1.07^3=R7350.26"

Interest earned amount:

"I=A-P=7350.26-6000=R1350.26"